OpenStudy (anonymous):

Lim of expression below as x approaches -4.

OpenStudy (anonymous):

\[\frac{ \sqrt{x^2-9}-5 }{ x+4 }\]

OpenStudy (tkhunny):

Why is that tricky? Denominator approaches zero numerator doesn't. This puppy blows up, no?

OpenStudy (anonymous):

man... i didn't even try to plug in 4 i just assumed i had to multiply by conjugate so i did and went on an endless expression... tthanks

OpenStudy (anonymous):

oh wait nevermind

OpenStudy (anonymous):

it is +9

OpenStudy (tkhunny):

Do NOT "plug in" anything, every. Give it some thought. Approaches zero. Approaches something finite. Absolutely do NOT substitute anything unless you know it is continuous. This isn't.

OpenStudy (tkhunny):

No, there is not a finite limit. It increases (or decreases) without bound.

OpenStudy (nfcfox):

Tkhunny you have to multiply by the conjugate.

OpenStudy (anonymous):

no i'm saying it should be x^2 +9

OpenStudy (anonymous):

typo

OpenStudy (nfcfox):

Let me type out the answer.

OpenStudy (tkhunny):

Now you don't. If you want more information, maybe that is so. To answer the problem, as stated, ain't nobody got time for that.

OpenStudy (anonymous):

ok so i did conjugate and got \[\frac{ x^2-16 }{ x \sqrt{x^2+9} +5x +4\sqrt{x^2+9}+20}\]

OpenStudy (anonymous):

tkhunny the expression i wrote had a typo, it is X^2+9 not - 9

OpenStudy (tkhunny):

Well, why didn't you say so?

OpenStudy (anonymous):

from there i tried L.H rule and got something and plugged in -4 and got -.2915 but that is wrong

OpenStudy (anonymous):

haha I tried

OpenStudy (nfcfox):

Let's start by multiplying the conjugate. We see that the root in the numerator as a - so we need to multiply by the conjugate. Let's start by multiplying the numerator by root(x^2 +9) +5 (This is the conjugate noticed how it is now +5). When we multiply the roots cancel and the 5 and -5 multiply to -25. We get: x^2 +9 -25= x^2 -16 Now the denominator: We can simply say (x+4)(root(x^2 +9) +5) But look! Our numerator can factor to (x-4) and (x+4). Now our x-4 in the numerator and denominator can cancel leaving us: x+4/(root(x^2 +9)+5) All we do now is simply plug in -4. We will get 0 in the numerator and 10 in the denom. however ultimately making the answer 0.

OpenStudy (nfcfox):

That's the answer. I am in calc but we haven't learned L. hop's rule but simply using the laws of limits at a constant we can get that.

OpenStudy (anonymous):

oh factor the numerator I see

OpenStudy (anonymous):

ok well i put 0 in for the answer and it is wrong

OpenStudy (nfcfox):

Yo i just saw my mistake

OpenStudy (nfcfox):

Can you re-enter it?

OpenStudy (anonymous):

lol no but i still want to figure it out.

OpenStudy (tkhunny):

Again, "plug in" doesn't mean anything. Never do it. What we have is this other expression, \(\dfrac{x-4}{\sqrt{x^{2}+9}+5}\), which is continuous at x = -4 and is equivalent to the original expression everywhere EXCEPT x = -4. Therefore, evaluating that last expression at x = -4, will provide the limit of the original expression as x approaches -4.

OpenStudy (nfcfox):

I canceled x-4 when the x+4 canceled making it -4-4 in the numerator which is -8. The denomater when plugging in -4 will =10 making the answer -4/5. Tell me if you need to see how I got 10 in the denominator.

OpenStudy (nfcfox):

No tkhunny. Stop. you do plug in after it is safe to. It's simply algebraic limits he doesn't need to look at continuity.

OpenStudy (tkhunny):

I dare you to find me a definition of "plug in". It doesn't exist.

OpenStudy (nfcfox):

Ugh tkhunny what have u taken?

OpenStudy (anonymous):

haha "Does not exist" get it

OpenStudy (nfcfox):

I mean are you in advanced mathematics? cause idk whether to argue or no

OpenStudy (anonymous):

well i now see how it is -4/5 thanks guys lolo

OpenStudy (tkhunny):

Let's just say that I learned a long time ago that careless language only confuses - never helps. There is a mathematical concept known as "substitution". There is no such thing called "plug in".